A projectile is thrown with speed 40 ms-1 at angle ? from horizontal. It is found that projectile as at same height at 1s and 3s. What is the angle of projection?

Grade:

2 Answers

Aastha
5 years ago
T=2usin¢/g :time of flight is given by this formula
Sin ¢ =gT/u
Now,
Sin¢=2*40/10=1/2
¢=30°
Putting the value of ¢in tan we get,
Tan 30°=1/√3
¢=tan-1(1/√3)
 
Khimraj
5 years ago
time period = 2*2 = 4 sec
so 2usin\Theta/g = 4
sin\Theta = 40/80 = ½
so \Theta = 30 degree
Hope it clears..........................